# Minimalistic and Scalable Quantum Reservoir Computing Enhanced with Feedback

### v1

@@ IV.2 Training quantum reservoir computing with feedback mechanism
The objective of training is to optimize the parameters $V_{n_{f}}$ and $W_{n}$ in order to achieve the actual output $y_{k}$ such that the normalized root mean square error
[…]
is minimized, where $L$ is the number of time steps in the training period and $\bar{y}_{k}$ represents the target output that captures some key features of the input.
To train the parameters $W_{n}$, a pseudoinverse method is sufficient in the regression analysis, since ${\rm NRMSE^{2}}$, being a quadratic function of $W_{n}$, only has a global minimum but no local minima. The readouts, $x_{kn}$, along with the constant bias term $x_{k0}=1$, are arranged in an $L\times\left(N_{readouts}+1\right)$ matrix ${\bf X}$. The target output, $\bar{y}_{k}$, is organized in an $L\times 1$ column vector ${\bf{\bf\bar{Y}}}$. The weights, $W_{n}$, are arranged in a $\left(N_{readouts}+1\right)\times 1$ column vector ${\bf W}$. The optimized weight ${\bf{\bf W}}$ that minimizes NRMSE is therefore determined by Dion2018 ; Govia2021

### v2

@@ IV.2 Training quantum reservoir computing with feedback mechanism
The objective of training is to optimize the parameters $V_{n_{f}}$ and $W_{n}$ in order to achieve the actual output $y_{k}$ such that the NRMSE defined in Eq. (9) is minimized.
To train the parameters $W_{n}$, a pseudoinverse method is sufficient in the regression analysis, since $\rm NRMSE^{2}$}, being a quadratic function of $W_{n}$, only has a global minimum but no local minima. The readouts, $x_{kn}$, along with the constant bias term $x_{k0}=1$, are arranged in an $L\times\left(N_{readouts}+1\right)$ matrix $\bf X$}. The target output, $\bar{y}_{k}$, is organized in an $L\times 1$ column vector $\bf{\bf\bar{Y}}$}. The weights, $W_{n}$, are arranged in a $\left(N_{readouts}+1\right)\times 1$ column vector $\bf W$}. The optimized weight $\bf{\bf W}$} that minimizes NRMSE is therefore determined by [Dion2018, Govia2021]
@@
Acknowledgements
We gratefully acknowledge the startup funding provided by Wyant College of Optical Sciences, University of Arizona, which supported the initial development of this research project.
Author Contributions
C.Z. and D.S. conceived the initial quantum reservoir computing system concept. C.Z, H.N., and P.E. contributed to the detailed operational design of the system. C.Z. conducted all numerical simulations. All authors contributed equally to the manuscript preparation.